OpenStudy (anonymous):

The following system is solved by substitution. Which expression is substituted for z in the second and third equations? x + 5y – z = 6 3x – y – 2z = 4 3x – 2y + 3z = 13

OpenStudy (anonymous):

Just isolate equations. So X+5Y-Z=6. Z=-6+X+5Y (Move 6 to the left side and Z to the right side). And then plug that into the 2nd equation. 3x-y-2z=4 so 3x-y-2(-6+x+5y)=4.